40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 291     Entry time: Mon Feb 6 11:19:13 2017
Author: Gabriele 
Type: Optics 
Category: Daily Progress 
Subject: Laser swap 

The Thorlabs laser has been misbehaving for the whole weekend. Even after many days being continuosly on, the wandering line is still moving all over the frequencies.

So this morning I swapped in a JDSU 1125P borrowed from the 40m lab, which provides about 6.8 mW of power. I tested it over the weekend on a separate test table, and after one day or so of operation the power looks reasonably stable. Now it's been on for a few hours: there is still a line moving around, but it's slowing down and hopefully setting down in a good place.

I started a series of test measurements on the samples that were already installed.

  • Quiet time before excitation: 1170443490
    Excitation (broad band) at 1170443523 (60 s)
    Quiet time after excitation: 1170443586

 

ELOG V3.1.3-